OpenStudy (anonymous):

-2(2x + 5) - 3 = -3(x - 1).

OpenStudy (anonymous):

-4x-10-3=-x+3 -4x+x=3+10+3 -3x=16 x=16/-3 x=-5.3

OpenStudy (anonymous):

not an option

OpenStudy (anonymous):

@ricky6511 would you still like help I can help you!